OpenStudy (anonymous):

A wife and husband, Linda and Mark, have ages that total 80 years. Two years ago, 2 times Marks age was 2 years more than, 3 times Linda's age. How old is each now?

OpenStudy (anonymous):

Linda is 32 and Mark is 48.

OpenStudy (radar):

Let x = Lindas age Ley y = Marks age x+y=80 yrs 2(y-2)=2+3(x-2) I think that was the way it was 2 yrs ago lol. 2y-4=2+3x-6 2y-3x=0 x+y=80 times 3 3y+3x=240 2y-3x=0 add 5y=240 y=48 x=32